# Integro-differential diffusion equations on graded Lie groups

### v1

@@ 3.2. Estimates in a homogeneous Sobolev space
Notice that in the last two theorems we were not able to gain any decay in the $L^{2}-$norm. This mainly happens since for the trivial representation on $\widehat{\mathbb{G}}$ the spectrum comes to 00. Now we recall a homogeneous Sololev space that involves the Rockland operator. In this space we can avoid somehow the contribution of the trivial representation and gain some time-decay in problem (1.1)-(1.2). For a detailed discussion on the Sobolev spaces in the setting of a graded Lie group, see [7, Subsection 4.4.1] and [8].
@@
Let $(k,l)\in\mathcal{PC}$ such that $l$ is positive and $l\in BV_{loc}(\mathbb{R}_{+}).$ Let $\mathbb{G}$ be a graded Lie group of homogeneous dimension $Q$. Let $\mathcal{R}$ be a positive Rockland operator of homogeneous degree $\nu$ on $\mathbb{G}$. For any $s\geqslant v$, if $u_{0}\in\dot{L}^{2}_{s-\nu}(\mathbb{G})$ then the solution of integro-differential difussion equation (1.1)-(1.2) $(f\equiv 0)$ belongs to $\dot{L}^{2}_{s}(\mathbb{G})$ and satisfies
@@
Let $(k,l)\in\mathcal{PC}$ such that $l$ is positive and $l\in BV_{loc}(\mathbb{R}_{+}).$ Let $\mathbb{G}$ be a graded Lie group of homogeneous dimension $Q$. Let $\mathcal{R}$ be a positive Rockland operator of homogeneous degree $\nu$ on $\mathbb{G}$. For any $s\geqslant v$, if $u_{0}\in\dot{L}^{2}_{s-\nu}(\mathbb{G})$ and $\bigg{(}\frac{l}{(1\ast l)(\cdot)}\ast f\bigg{)}(t)\in\dot{L}^{2}_{s-\nu}(\mathbb{G})$ then the solution of the integro-differential difussion equation (1.1)-(1.2) belongs to $\dot{L}^{2}_{s}(\mathbb{G})$ and satisfies
@@
$\begin{split}\partial_{t}\left(k\ast\left[\widehat{u}(t,\pi)_{ij}-\widehat{u_{0}}(\pi)_{ij}\right]\right)+\pi_{i}^{2}\widehat{u}(t,\pi)_{ij}&=\widehat{f}(t,\pi)_{ij},\quad t>0,\\ \widehat{u}(t,\pi)_{ij}|_{{}_{{}_{t=0}}}&=\widehat{u_{0}}(\pi)_{ij},\end{split}$
where we are considering any $i,j\in\mathbb{N}$ and any $\pi\in\widehat{\mathbb{G}}.$ By [21, Lemma 5.1 and Lemma 5.3], the solution is given by
@@
where $\overline{r_{ij}}_{\pi_{i}^{2}}$ is a nonnegative and nonincreasing function. Moreover, by estimate (2.3) and [21, Formula (5.9)] we arrive at
@@ 4.1. Euclidean space
$(x,y,t)(x^{\prime},y,t^{\prime}):=(x+x^{\prime},y+y^{\prime},t+t^{\prime}+\frac{1}{2}(xy^{\prime}-x^{\prime}y))\,,$
where $(x,y,t),(x^{\prime},y,t^{\prime})\in\mathbb{H}^{n}$. The canonical basis of its Lie algebra is given by the (left-invariant) vector fields
@@ 5.1. Time-fractional diffusion equations
Then the above $L^{p}(\mathbb{G})\,-\,L^{q}(\mathbb{G})$ estimate coincides with the estimate obtained in [10, Theorem 3.3] for the problem (5.1).
@@ 5.4. Multi-term time-fractional diffusion equation
In [17, Lemma 3.2] it is proven that if $\mu\leqslant|\arg(z_{1})|\leqslant\pi,\,\alpha_{1}\pi/2<\mu<\alpha_{1}\pi$ and there exists $K>0$ such that $-K\leqslant z_{j}<0\,\,(j=2,...,m)$, then there exists a constant $C>0$ depending only on $\mu,$ $K$, $\alpha_{j}\,(j=1,...,m)$ and $\beta$ such that
@@
This result improves decay estimates from [2, Subsection 4.3] and [24], where decay estimates of local in time solutions were obtained with the time-dependent constants.
@@ 6. Equations on compact Lie groups
Let $(k,l)\in\mathcal{PC}$ such that $l$ is positive and $l\in BV_{loc}(\mathbb{R}_{+}).$ Let $\mathrm{G}$ be a compact Lie group. Let $\mathcal{L}$ be a positive linear left invariant operator on $\mathrm{G}$ (maybe unbounded) such that
@@
Let $(k,l)\in\mathcal{PC}$ such that $l$ is positive and $l\in BV_{loc}(\mathbb{R}_{+}).$ Let $\mathrm{G}$ be a compact Lie group. Let $\mathcal{L}$ be a positive linear left invariant operator on $\mathrm{G}$ (maybe unbounded) such that condition (6.2) is true. If $u_{0}$ is a continuous function in $L^{p}(\mathrm{G})$ $(1<p<+\infty)$, $f\in C([0,T];L^{p}(\mathrm{G}))$ and $l\ast f\in C([0,T];L^{p}(\mathrm{G}))$ then the integral equation (6.1) is well-posed. Also, if the condition (6.3) holds, $u_{0}\in L^{p}(\mathrm{G})$ $(1<p\leqslant 2)$ and $l\ast f\in W^{1,1}([0,T];L^{p}(\mathrm{G}))$ then the following mild solution of (6.1)

### v2

@@ 3.2. Estimates in a homogeneous Sobolev space
Notice that in the last two theorems we were not able to gain any decay in the $L^{2}-$norm. This mainly happens since for the trivial representation on $\widehat{\mathbb{G}}$ the spectrum comes to 00. Now we recall a homogeneous Sololev space that involves the Rockland operator. In this space we can avoid somehow the contribution of the trivial representation and gain some time-decay in problem (1.1)-(1.2). For a detailed discussion on the Sobolev spaces in the setting of a graded Lie group, see [8, Subsection 4.4.1] and [9].
@@
Let $(k,l)\in\mathcal{PC}$ be such that $l$ is positive and $l\in BV_{loc}(\mathbb{R}_{+}).$ Let $\mathbb{G}$ be a graded Lie group of homogeneous dimension $Q$. Let $\mathcal{R}$ be a positive Rockland operator of homogeneous degree $\nu$ on $\mathbb{G}$. For any $s\geqslant v$, if $u_{0}\in\dot{L}^{2}_{s-\nu}(\mathbb{G})$ then the solution of integro-differential difussion equation (1.1)-(1.2) $(f\equiv 0)$ belongs to $\dot{L}^{2}_{s}(\mathbb{G})$ and satisfies
@@
Let $(k,l)\in\mathcal{PC}$ be such that $l$ is positive and $l\in BV_{loc}(\mathbb{R}_{+}).$ Let $\mathbb{G}$ be a graded Lie group of homogeneous dimension $Q$. Let $\mathcal{R}$ be a positive Rockland operator of homogeneous degree $\nu$ on $\mathbb{G}$. For any $s\geqslant v$, if $u_{0}\in\dot{L}^{2}_{s-\nu}(\mathbb{G})$ and $\bigg{(}\frac{l}{(1\ast l)(\cdot)}\ast f\bigg{)}(t)\in\dot{L}^{2}_{s-\nu}(\mathbb{G})$ then the solution of the integro-differential difussion equation (1.1)-(1.2) belongs to $\dot{L}^{2}_{s}(\mathbb{G})$ and satisfies
@@
$\begin{split}\partial_{t}\left(k\ast\left[\widehat{u}(t,\pi)_{ij}-\widehat{u_{0}}(\pi)_{ij}\right]\right)+\pi_{i}^{2}\widehat{u}(t,\pi)_{ij}&=\widehat{f}(t,\pi)_{ij},\quad\widehat{u}(t,\pi)_{ij}|_{{}_{{}_{t=0}}}=\widehat{u_{0}}(\pi)_{ij};\quad t>0,\end{split}$
where we are considering any $i,j\in\mathbb{N}$ and any $\pi\in\widehat{\mathbb{G}}.$ By [28, Lemma 5.1 and Lemma 5.3], the solution is given by
@@
where $\overline{r_{ij}}_{\pi_{i}^{2}}$ is a nonnegative and nonincreasing function. Moreover, by estimate (2.3) and [28, Formula (5.9)] we arrive at
@@ 4.1. Euclidean space
$(x,y,t)(x^{\prime},y^{\prime},t^{\prime}):=(x+x^{\prime},y+y^{\prime},t+t^{\prime}+\frac{1}{2}(xy^{\prime}-x^{\prime}y))\,,$
where $(x,y,t),(x^{\prime},y^{\prime},t^{\prime})\in\mathbb{H}^{n}$. The canonical basis of its Lie algebra is given by the (left-invariant) vector fields
@@ 5.1. Time-fractional diffusion equations
Then the above $L^{p}(\mathbb{G})\,-\,L^{q}(\mathbb{G})$ estimate coincides with the estimate obtained in [11, Theorem 3.3] for the problem (5.1).
@@ 5.4. Multi-term time-fractional diffusion equation
In [19, Lemma 3.2] it is proven that if $\mu\leqslant|\arg(z_{1})|\leqslant\pi,\,\alpha_{1}\pi/2<\mu<\alpha_{1}\pi$ and there exists $K>0$ such that $-K\leqslant z_{j}<0\,\,(j=2,...,m)$, then there exists a constant $C>0$ depending only on $\mu,$ $K$, $\alpha_{j}\,(j=1,...,m)$ and $\beta$ such that
@@
This result improves decay estimates from [32], where decay estimates of local in time solutions were obtained with the time-dependent constants.
@@ 6. Equations on compact Lie groups
Let $(k,l)\in\mathcal{PC}$ be such that $l$ is positive and $l\in BV_{loc}(\mathbb{R}_{+}).$ Let $\mathrm{G}$ be a compact Lie group. Let $\mathcal{L}$ be a positive linear left invariant operator on $\mathrm{G}$ (maybe unbounded) such that
@@
Let $(k,l)\in\mathcal{PC}$ be such that $l$ is positive and $l\in BV_{loc}(\mathbb{R}_{+}).$ Let $\mathrm{G}$ be a compact Lie group. Let $\mathcal{L}$ be a positive linear left invariant operator on $\mathrm{G}$ (maybe unbounded) such that condition (6.2) is true. If $u_{0}$ is a continuous function in $L^{p}(\mathrm{G})$ $(1<p<+\infty)$, $f\in C([0,T];L^{p}(\mathrm{G}))$ and $l\ast f\in C([0,T];L^{p}(\mathrm{G}))$ then the integral equation (6.1) is well-posed. Also, if the condition (6.3) holds, $u_{0}\in L^{p}(\mathrm{G})$ $(1<p\leqslant 2)$ and $l\ast f\in W^{1,1}([0,T];L^{p}(\mathrm{G}))$ then the following mild solution of (6.1)